# Smart Hiring Redefined: An Intelligent Recruitment Management Platform

### v2

@@ 1 INTRODUCTION
At present, online recruitment platforms that provide digital hiring and presentation services function as crucial channels of interaction between enterprises and potential candidates. Many universities respond to this trend by implementing similar systems to disseminate employment information and coordinate recruitment activities. Prominent platforms such as Lagou and Kanzhun Limited, which gain increasing visibility in recent years, position themselves as providers of intelligent and personalized recruitment services [33]. Local governments likewise launch campus recruitment platforms that focus on strengthening collaboration between regional businesses and universities and encourage graduates to pursue employment locally. Internationally, platforms such as LinkedIn offer comprehensive recruitment services that reinforce ties between universities and enterprises, while emerging startups leverage artificial intelligence and big data technologies to deliver more customized services. Overseas campus recruitment systems further emphasize career guidance and personalized recommendations, underscoring the global shift toward more data driven and candidate centric approaches[37].
Despite the rapid diffusion of intelligent campus recruitment systems worldwide, significant disparities persist with respect to their technological sophistication, service quality, and level of intelligence. Broadly, system functionalities fall into two principal categories: recruitment and presentation[32]. The recruitment function enables enterprises to publish job postings, allowing students to browse, select, and apply for desired positions online and await responses from employers. The presentation function permits enterprises to conduct on campus sessions and disseminate presentation information, subject to administrative approval, while allowing students to decide independently whether to attend[4].
From an organizational perspective, these systems enable human resource professionals to more efficiently evaluate the skills and potential of graduating students, thereby offering faster and more cost effective recruitment channels and reducing advertising and promotional expenditures. From a student perspective, the systems facilitate timely access to career related information and industry insights, minimize the time cost associated with online searches, and enhance early career planning. Participation in recruitment presentations further increases opportunities for direct engagement with enterprises and supports informed career decision making[18][39].
@@ 2.1.1 Technical Feasibility
This system adopts a B/S architecture and is designed based on the J2EE three tier structure (presentation layer, business logic layer, and data service layer). It utilizes Spring Boot to integrate the SSH framework (Spring 4, Struts 2, Hibernate 4) as its technical foundation, combined with a MySQL relational database. This ensures the system’s efficiency, stability, and maintainability. Considering the existing Java technology environment, the implementation of this system is technically feasible[38].
@@ 2.1.3 Social Feasibility
In today’s fiercely competitive job market, online recruitment emerges as an increasingly vital component of campus hiring. With no geographical constraints, broad reach, and extended duration, it enables rapid dissemination of job openings and attracts a large pool of applicants. As the internet continues to evolve rapidly, this recruitment method becomes mainstream, making it a viable approach from a societal perspective.
@@ 2.2.1 Administrator Use Case
As the core administrators of the system, administrators bear primary responsibilities including identity authentication, corporate information maintenance, and student information management. After accessing the system through secure login verification, administrators review corporate registration applications and perform operations to add, delete, modify, and query corporate information, ensuring the integrity and accuracy of corporate data[31]. Simultaneously, administrators possess comprehensive permissions to maintain student basic information, supporting the ability to query, modify, and delete student records by student ID. The system design prioritizes operational security and fault tolerance, incorporating reset functions, confirmation prompts, and exception handling mechanisms. These features effectively prevent operational errors and safeguard data consistency, thereby maintaining the orderly operation of the recruitment platform[30][21].
@@ 2.3 Students Use Case
The Student Use Case Diagram clearly illustrates the comprehensive job-seeking service system for student users in the system. With “Student” as the central actor, the diagram integrates three core functionalities within the Recruitment Management use case module. Students access real time job postings through the “View Recruitment Information” feature, submit applications via the “Submit Resume” function, and continuously track application progress using the “View Application Status” capability[7][19].
@@ 2.4 Bussiness Use Case
The Enterprise Business Management Use Case Diagram comprehensively illustrates the full featured management system for corporate users. Centered on the “Enterprise” as the core actor, the diagram presents a complete solution that encompasses information management, recruitment operations, and talent interaction. Corporate users utilize the “Enterprise Information Management” use case module to add, view, modify, and delete basic information, thereby enabling comprehensive maintenance of their profiles[15][35].
@@ 3 System Architecture
This system adopts a classic layered architecture pattern, dividing system functionality into four logical layers to ensure separation of responsibilities and modular development. The foundation consists of an infrastructure layer and a data layer[23]. The infrastructure layer provides the system’s fundamental operational environment, including networking, computing, and storage, while the data layer enables persistent data storage and management through a relational database[16].
The middle layer comprises the service layer and the application layer. The service layer encapsulates reusable core business components, offering unified data access and service interfaces to upper layers. The application layer constructs specific functional modules tailored to different user roles and business scenarios[41]. The top layer is the presentation layer, responsible for rendering and displaying the user interface. Communication between layers occurs through well defined interfaces, with lower layers providing service support to upper layers. This establishes strict dependency relationships, enabling the system to achieve high cohesion and low coupling[24].
@@ 4 IMPLEMENTATION
This login method employs an MVC layered architecture and performs authentication through the service layer. It uses an HQL query to validate administrator credentials (See Listing 1 for Details). Upon successful matching, the system stores the administrator object in the session and returns a success token; otherwise, it returns an error message. This implementation demonstrates the separation of business logic from data access[14].
@@ 14 String hql = "FROM Admin WHERE adminId = ? AND adminPassword = ?";
This method retrieves the complete list of presentation applications and passes the data to the frontend view via ActionContext(See Listing 2 for Details). Upon successful execution, it returns the string success, which triggers the corresponding view rendering to complete frontend–backend data interaction. The code embodies the fundamental functions of a controller within the MVC architecture.
@@
This method processes student password modification requests within the MVC based system architecture. The controller layer first verifies the consistency between the new password and its confirmation to ensure input validity. It then invokes the business logic layer to execute the password update operation under controlled transactional conditions(See Listing 3 for Details). The system implements input validation, exception handling, and secure password storage to enhance data integrity and operational security. Based on the result returned by the business layer, the controller returns the corresponding success or failure view state, thereby reinforcing the separation of concerns and maintainability within the layered architecture[32].
@@
This method enables students to query their scheduled presentation arrangements within the system. The service layer first verifies the student’s identity to ensure secure and authorized access(See Listing 4 for Details). It then retrieves all scheduled activities associated with the student and stores the resulting data in the request context. Finally, the controller returns a success status to trigger the appropriate view rendering, thereby completing the full process from data retrieval to presentation. This design reinforces secure data access, preserves data integrity, and illustrates the separation of concerns within the layered architecture.
@@
This controller class handles requests related to presentation activities. It invokes the findAll() method in the service layer to retrieve the complete list of activities and stores the results in the arclist property(See Listing 6 for Details). Upon returning a success status, the corresponding view renders the data for display. This design demonstrates the loose coupling between the controller and the view within the MVC architecture, thereby enhancing maintainability and scalability.
@@ 6 CONCLUTION
The design and implementation of this system establish an intelligent recruitment management platform based on a B/S architecture and the J2EE three tier framework. By integrating the SpringBoot framework with SSH components and a MySQL database, the system demonstrates robust stability, security, and maintainability. Test results confirm that the three core modules student interface, enterprise interface, and administrator interface achieve complete functional implementation. Business processes operate smoothly, and user interactions align with expectations, thereby validating the rationality of the system design and the feasibility of the technical solution. Although the system achieves its fundamental design objectives, opportunities for improvement remain in areas such as high concurrency processing, interface responsiveness optimization, and mobile device adaptation. Future enhancements can incorporate caching mechanisms and asynchronous processing to boost system performance, while expanding data analytics capabilities to elevate the intelligence of recruitment processes, thereby better serving the campus recruitment ecosystem.

### v4

@@ 1 INTRODUCTION
At present, online recruitment platforms that provide digital hiring and presentation services function as crucial channels of interaction between enterprises and potential candidates. Many universities respond to this trend by implementing similar systems to disseminate employment information and coordinate recruitment activities. Prominent platforms such as Lagou and Kanzhun Limited, which gain increasing visibility in recent years, position themselves as providers of intelligent and personalized recruitment services [52]. Local governments likewise launch campus recruitment platforms that focus on strengthening collaboration between regional businesses and universities and encourage graduates to pursue employment locally. Internationally, platforms such as LinkedIn offer comprehensive recruitment services that reinforce ties between universities and enterprises, while emerging startups leverage artificial intelligence and big data technologies to deliver more customized services. Overseas campus recruitment systems further emphasize career guidance and personalized recommendations, underscoring the global shift toward more data driven and candidate centric approaches[57][38].
Despite the rapid diffusion of intelligent campus recruitment systems worldwide, significant disparities persist with respect to their technological sophistication, service quality, and level of intelligence. Broadly, system functionalities fall into two principal categories: recruitment and presentation[51]. The recruitment function enables enterprises to publish job postings, allowing students to browse, select, and apply for desired positions online and await responses from employers. The presentation function permits enterprises to conduct on campus sessions and disseminate presentation information, subject to administrative approval, while allowing students to decide independently whether to attend[10][33].
From an organizational perspective, these systems enable human resource professionals to more efficiently evaluate the skills and potential of graduating students, thereby offering faster and more cost effective recruitment channels and reducing advertising and promotional expenditures[12]. From a student perspective, the systems facilitate timely access to career related information and industry insights, minimize the time cost associated with online searches, and enhance early career planning. Participation in recruitment presentations further increases opportunities for direct engagement with enterprises and supports informed career decision making[29][59][3].
@@ 2.1.1 Technical Feasibility
This system adopts a B/S architecture and is designed based on the J2EE three tier structure (presentation layer, business logic layer, and data service layer). It utilizes Spring Boot to integrate the SSH framework (Spring 4, Struts 2, Hibernate 4) as its technical foundation, combined with a MySQL relational database. This ensures the system’s efficiency, stability, and maintainability[56]. Considering the existing Java technology environment, the implementation of this system is technically feasible[58][35].
@@ 2.1.3 Social Feasibility
In today’s fiercely competitive job market, online recruitment emerges as an increasingly vital component of campus hiring. With no geographical constraints, broad reach, and extended duration, it enables rapid dissemination of job openings and attracts a large pool of applicants[20]. As the internet continues to evolve rapidly, this recruitment method becomes mainstream, making it a viable approach from a societal perspective[18].
@@ 2.2.1 Administrator Use Case
As the core administrators of the system, administrators bear primary responsibilities including identity authentication, corporate information maintenance, and student information management. After accessing the system through secure login verification, administrators review corporate registration applications and perform operations to add, delete, modify, and query corporate information, ensuring the integrity and accuracy of corporate data[50]. Simultaneously, administrators possess comprehensive permissions to maintain student basic information, supporting the ability to query, modify, and delete student records by student ID. The system design prioritizes operational security and fault tolerance, incorporating reset functions, confirmation prompts, and exception handling mechanisms[46]. These features effectively prevent operational errors and safeguard data consistency, thereby maintaining the orderly operation of the recruitment platform[49][32][5].
@@ 2.3 Students Use Case
The Student Use Case Diagram clearly illustrates the comprehensive job-seeking service system for student users in the system. With “Student” as the central actor, the diagram integrates three core functionalities within the Recruitment Management use case module. Students access real time job postings through the “View Recruitment Information” feature, submit applications via the “Submit Resume” function, and continuously track application progress using the “View Application Status” capability[15][30][46].
@@ 2.4 Bussiness Use Case
The Enterprise Business Management Use Case Diagram comprehensively illustrates the full featured management system for corporate users[11]. Centered on the “Enterprise” as the core actor, the diagram presents a complete solution that encompasses information management, recruitment operations, and talent interaction. Corporate users utilize the “Enterprise Information Management” use case module to add, view, modify, and delete basic information, thereby enabling comprehensive maintenance of their profiles[26][54].
@@ 3 System Architecture
This system adopts a classic layered architecture pattern, dividing system functionality into four logical layers to ensure separation of responsibilities and modular development. The foundation consists of an infrastructure layer and a data layer[40]. The infrastructure layer provides the system’s fundamental operational environment, including networking, computing, and storage, while the data layer enables persistent data storage and management through a relational database[27][1].
The middle layer comprises the service layer and the application layer. The service layer encapsulates reusable core business components, offering unified data access and service interfaces to upper layers. The application layer constructs specific functional modules tailored to different user roles and business scenarios[61][2]. The top layer is the presentation layer, responsible for rendering and displaying the user interface. Communication between layers occurs through well defined interfaces, with lower layers providing service support to upper layers. This establishes strict dependency relationships, enabling the system to achieve high cohesion and low coupling[41].
@@ 4 IMPLEMENTATION
This login method employs an MVC layered architecture and performs authentication through the service layer. It uses an HQL query to validate administrator credentials (See Listing 1 for Details). Upon successful matching, the system stores the administrator object in the session and returns a success token; otherwise, it returns an error message. This implementation demonstrates the separation of business logic from data access[24][34].
@@ 14 String hql = "FROM Admin WHERE adminId = ? AND adminPassword = ?";
This method retrieves the complete list of presentation applications and passes the data to the frontend view via ActionContext(See Listing 2 for Details). Upon successful execution, it returns the string success, which triggers the corresponding view rendering to complete frontend–backend data interaction. The code embodies the fundamental functions of a controller within the MVC architecture[36].
@@
This method processes student password modification requests within the MVC based system architecture. The controller layer first verifies the consistency between the new password and its confirmation to ensure input validity. It then invokes the business logic layer to execute the password update operation under controlled transactional conditions(See Listing 3 for Details). The system implements input validation, exception handling, and secure password storage to enhance data integrity and operational security. Based on the result returned by the business layer, the controller returns the corresponding success or failure view state, thereby reinforcing the separation of concerns and maintainability within the layered architecture[51][42].
@@
This method enables students to query their scheduled presentation arrangements within the system. The service layer first verifies the student’s identity to ensure secure and authorized access(See Listing 4 for Details). It then retrieves all scheduled activities associated with the student and stores the resulting data in the request context. Finally, the controller returns a success status to trigger the appropriate view rendering, thereby completing the full process from data retrieval to presentation. This design reinforces secure data access, preserves data integrity, and illustrates the separation of concerns within the layered architecture[25][4].
@@
This controller class handles requests related to presentation activities. It invokes the findAll() method in the service layer to retrieve the complete list of activities and stores the results in the arclist property(See Listing 6 for Details). Upon returning a success status, the corresponding view renders the data for display. This design demonstrates the loose coupling between the controller and the view within the MVC architecture, thereby enhancing maintainability and scalability[8].
@@ 6 CONCLUTION
The design and implementation of this system establish an intelligent recruitment management platform based on a B/S architecture and the J2EE three tier framework. By integrating the SpringBoot framework with SSH components and a MySQL database, the system demonstrates robust stability, security, and maintainability. Test results confirm that the three core modules student interface, enterprise interface, and administrator interface achieve complete functional implementation. Business processes operate smoothly, and user interactions align with expectations, thereby validating the rationality of the system design and the feasibility of the technical solution. Although the system achieves its fundamental design objectives, opportunities for improvement remain in areas such as high concurrency processing, interface responsiveness optimization, and mobile device adaptation. Future enhancements can incorporate caching mechanisms and asynchronous processing to boost system performance, while expanding data analytics capabilities to elevate the intelligence of recruitment processes, thereby better serving the campus recruitment ecosystem[37].